# Recent Advances in Large Langauge Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation

### v1

@@ 4 Dynamic Benchmarking
Due to the inherent limitations of static benchmarking schemes, they face challenges in providing a transparent yet faithful evaluation of LLMs. To address this, dynamic benchmarking has been proposed.
@@ 4.1 Problem Formulation
Dynamic Mechnisms Bechmark Name Evaluation Creteria
Correctness Scalability Collision Stable of Complexity Diversity Interpretability
@@
Self-Evolving (Wang et al., 2024a)
@@
4.2 Evaluation Criteria
While many dynamic benchmarking methods have been proposed to evaluate LLMs, the evaluation criteria for assessing these benchmarks themselves remain non-standardized. To this end, we propose the following evaluation criteria to assess the quality of a dynamic benchmarking algorithm.
@@ 4.2.1 Correctness
The first criterion for evaluating the quality of dynamic benchmarking is Correctness. If the correctness of the generated dataset cannot be guaranteed, the benchmark may provide a false sense of reliability when applied to benchmarking LLMs, leading to misleading evaluations. We quantify the correctness of dynamic benchmarks as:
@@
where $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$ represent the input and output of the $i^{th}$ transformation, respectively. The function $\mathcal{G}(\cdot)$ is an oracle that returns the ground truth of its input, ensuring an objective reference for correctness evaluation. For example, the function $\mathcal{G}(\cdot)$ could be a domain-specific an annotator. This equation can be interpreted as the expected alignment between the transformed dataset’s outputs and their corresponding ground truth values, measured using the scoring function $\mathcal{S}(\cdot)$. A higher correctness score indicates that the dynamic benchmark maintains correctness to the ground truth.
@@ 4.2.2 Scalability
The next evaluation criterion is scalability, which measures the ability of dynamic benchmarking methods to generate large-scale benchmark datasets. A smaller dataset can introduce more statistical errors during the benchmarking process. Therefore, an optimal dynamic benchmark should generate a larger dataset while minimizing associated costs. The scalability of a dynamic benchmark is quantified as:
@@
This represents the expectation over the entire transformation space, where $\lVert T_{i}(\mathcal{D})\rVert$ is the size of the transformed dataset, and $\lVert\mathcal{D}\rVert$ is the size of the original dataset. The function $\textsf{Cost}(\cdot)$ measures the cost associated with the transformation process, which could include monetary cost, time spent, or manual effort according to the detailed scenarios. This equation could be interpreted as the proportion of data that can be generated per unit cost.
@@ 4.2.3 Collision
One of the main motivations for dynamic benchmarking is to address the challenge of balancing transparent benchmarking with the risk of data contamination. Since the benchmarking algorithm is publicly available, an important concern arises: If these benchmarks are used to train LLM, can they still reliably reflect the true capabilities of LLMs? To evaluate the robustness of a dynamic benchmark against this challenge, we introduce the concept of collision in dynamic benchmarking. Collision refers to the extent to which different transformations of the benchmark dataset produce overlapping data, potentially limiting the benchmark’s ability to generate novel and diverse test cases. To quantify this, we propose the following metrics:
Collision $\begin{split}&\textsf{Collision Rate}=\mathbb{E}_{\begin{subarray}{c}i,j=1,\,i\neq j\end{subarray}}^{N}\left[\frac{\lVert\mathcal{D}_{i}\cap\mathcal{D}_{j}\rVert}{\lVert\mathcal{D}\rVert}\right]\\ &\textsf{Repeat}=\mathbb{E}_{i=1}^{N}\left[k\mid k=\min\left\{\bigcup_{j=1}^{k}\mathcal{D}_{j}\supseteq\mathcal{D}_{i}\right\}\right]\end{split}$
Collision Rate measures the percentage of overlap between two independently transformed versions of the benchmark dataset, indicating how much poential contamination among two trials. Repeat Trials quantifies the expected number of transformation trials required to fully regenerate an existing transformed dataset $T_{i}(\mathcal{D})$, providing insight into the benchmark’s ability to produce novel variations. These metrics help assess whether a dynamic benchmark remains effective in evaluating LLM capabilities, even when exposed to potential training data contamination.
4.2.4 Stable of Complexity
Dynamic benchmarks must also account for complexity to help users determine whether a performance drop in an LLM on the transformed dataset is due to potential data contamination or an increase in task complexity. If a dynamic transformation increases the complexity of the seed dataset, a performance drop is expected, even without data contamination. However, accurately measuring the complexity of a benchmark dataset remains a challenging task. Existing work has proposed various complexity metrics, but these are often domain-specific and do not generalize well across different applications. For example, DyVal (Zhu et al., 2024a) proposes applying graph complexity to evaluate the complexity of reasoning problems. Formally, given a complexity measurement function $\Psi(\cdot)$, the stability can be formulated as:
@@ 4.2.5 Diversity
Besides the aforementioned criteria, another important factor is the diversity of the transformed dataset. This diversity can be categorized into two components: external diversity and internal diversity: External diversity measures the variation between the transformed dataset and the seed dataset. Internal diversity quantifies the differences between two transformation trials.
@@
where $\Theta(\cdot)$ is a function that measures the diversity between two datasets. For example, it could be the N-gram metrics or the reference based metrics, such as BLEU scores.
@@ 4.2.6 Interpretability
Dynamic benchmarking generates large volumes of transformed data, making manual verification costly and challenging. To ensure correctness, the transformation process must be interpretable. Interpretable transformations reduce the need for extensive manual validation, lowering costs. Rule-based or manually crafted transformations are inherently interpretable, while LLM-assisted transformations depend on the model’s transparency and traceability. In such cases, additional mechanisms like explainability tools, or human-in-the-loop validation may be needed to ensure reliability and correctness.
@@ 4.3 Existing Work
Building on the task formats of static benchmarks, dynamic benchmarks have been introduced to assess LLM capabilities while minimizing data contamination and ensuring fairness. Table 2 summarizes recent dynamic benchmarks for LLM evaluation. Dynamic benchmarks can be categorized into four types based on their construction process: temporal cutoff, rule-based generation, LLM-based generation, and hybrid approaches. Temporal cutoff follows a data collection process similar to static benchmarks, with the key difference being that data is gathered from newly released information. Rule-based and LLM-based generation approaches create novel evaluation data points using predefined rules or by leveraging the strong generative capabilities of LLMs. Finally, hybrid approaches combine the idea of these different approaches.
@@ 4.3.1 Temporal Cutoff
Since LLMs typically have a knowledge cutoff date, using data collected after this cutoff to construct dataset can help evaluate the model while mitigating data contamination. This approach has been widely adopted to construct reliable benchmarks that prevent contamination. LiveBench (White et al., 2024) collects questions based on the latest information source, e.g., math competitions from the past 12 months, with new questions added and updated every few months. AntiLeak-Bench (Wu et al., 2024) generates queries about newly emerged knowledge that was unknown before the model’s knowledge cutoff date to eliminate potential data contamination. AcademicEval (Zhang et al., 2024a) designs academic writing tasks on latest arXiv papers. LiveCodeBench (Jain et al., 2024) continuously collects new human-written coding problems from online coding competition platforms like LeetCode. LiveAoPSBench (Mahdavi et al., 2025) collects live math problems from the Art of Problem Solving forum. Forecastbench (Karger et al., 2024) updates new forecasting questions on a daily basis from different data sources, e.g., prediction markets.
@@
The collection process typically requires significant human effort (White et al., 2024; Jain et al., 2024), and continuous updates demand ongoing human involvement. Despite the popularity of temporal cutoffs, using recent information from competitions to evaluate LLMs can still lead to data contamination, as these problems are likely to be reused in future competitions (Wu et al., 2024). Verification is often overlooked in these live benchmarks (White et al., 2024).
4.3.2 Rule-Based Generation
This method synthesizes new test cases based on predefined rules, featuring an extremely low collision probability (Zhu et al., 2024a).
Template-Based
GSM-Symbolic (Mirzadeh et al., 2025) creates dynamic math benchmarks by using query templates with placeholder variables, which are randomly filled to generate diverse problem instances. Mathador-LM(Kurtic et al., 2024) generates evaluation queries by adhering to the rules of Mathador games(Puma et al., 2023) and varying input numbers. MMLU-CF (Zhao et al., 2024) follows the template of multiple-choice questions and generates novel samples by shuffling answer choices and randomly replacing incorrect options with "None of the other choices."
Table-Based
S3Eval (Lei et al., 2024) evaluates the reasoning ability of LLMs by assessing their accuracy in executing random SQL queries on randomly generated SQL tables.
Graph-Based
In this category, LLMs are evaluated with randomly generated graphs. For instance, DyVal (Zhu et al., 2024a) assesses the reasoning capabilities of LLMs using randomly generated directed acyclic graphs (DAGs). The framework first constructs DAGs with varying numbers of nodes and edges to control task difficulty. These DAGs are then transformed into natural language descriptions through rule-based conversion. Finally, the LLM is evaluated by querying it for the value of the root node. Similarly, NPHardEval (Fan et al., 2024) evaluates the reasoning ability of LLMs on well-known P and NP problems, such as the Traveling Salesman Problem (TSP). Random graphs of varying sizes are synthesized as inputs for TSP to assess the LLM’s performance. Xie et al. (2024) automatically constructs Knights and Knaves puzzles with random reasoning graph.
Limitations
The pre-defined rules may limit sample diversity, and publicly available rule-generated data may increase the risk of in-distribution contamination during training (Tu et al., 2024).
4.3.3 LLM-Based Generation
Benchmark Rewriting
In this category, LLMs are employed to rewrite samples from existing static benchmarks, which may be contaminated. Auto-Dataset (Ying et al., 2024) prompts LLMs to generate two types of new samples: one that retains the stylistics and essential knowledge of the original, and another that presents related questions at different cognitive levels (Bloom et al., 1956). StructEval (Cao et al., 2024) expands on examined concepts from the original benchmark by using LLMs and knowledge graphs to develop a series of extended questions. ITD (Zhu et al., 2024c) utilizes a contamination detector (Shi et al., 2024) to identify contaminated samples in static benchmarks and then prompts an LLM to rewrite them while preserving their difficulty levels. VarBench (Qian et al., 2024) prompts LLMs to identify and replace variables in samples from existing benchmarks, generating new samples.
Interactive Evaluation
In this category, inspired by the human interview process, LLMs are evaluated through multi-round interactions with an LLM  (Li et al., 2023b). LLM-as-an-Interviewer (Kim et al., 2024) employs an interviewer LLM that first paraphrases queries from existing static benchmarks and then conducts a multi-turn evaluation by posing follow-up questions or providing feedback on the examined LLM’s responses. TreeEval (Li et al., 2024b) begins by generating an initial question on a given topic using an LLM. Based on the previous topic and the examined LLM’s response, it then generates follow-up subtopics and corresponding questions to further assess the model. KIEval (Yu et al., 2024) generates follow-up questions based on the evaluated model’s response to an initial question from a static benchmark.
Multi-Agent Evaluation
Inspired by the recent success of multi-agents systems (Guo et al., 2024), multi-agent collaborations are used to construct dynamic benchmarks. Benchmark Self-Evolving (Wang et al., 2024a) employs a multi-agent framework to dynamically extend existing static benchmarks, showcasing the potential of agent-based methods. Given a task description, BENCHAGENTS (Butt et al., 2024) leverages a multi-agent framework for automated benchmark creation. It splits the process into planning, generation, verification, and evaluation—each handled by a specialized LLM agent. This coordinated approach, with human-in-the-loop feedback, yields scalable, diverse, and high-quality benchmarks.
Limitations
The quality of LLM-generated samples is often uncertain. For example, human annotation of LatestEval (Li et al., 2023d) shows that about 10% of samples suffer from faithfulness and answerability issues, compromising evaluation reliability. Moreover, in interactive evaluation, reliability largely depends on the interviewer LLM, which generates questions and assigns scores.
4.3.4 Hybrid Generation
LatestEval (Li et al., 2023d) combines temporal cutoff and LLM-based generation to automatically generate reading comprehension datasets using LLMs on real-time content from sources such as BBC. DARG (Zhang et al., 2024b) integrates LLM-based and graph-based generation. It first extracts reasoning graphs from existing benchmarks and then perturbs them into new samples using predefined rules. C2LEVA (Li et al., 2024c) incorporates all three contamination-free construction methods to build a contamination-free bilingual evaluation.
5 Discussions
Key Insights. Our analysis highlights two key insights regarding data contamination in LLM benchmarking. First, static benchmarks become less effective as training corpora grow. The probability of contamination increases with $\text{Pr}_{\text{contam}}\propto|\mathcal{D}_{\text{train}}|\cdot|\mathcal{D}_{\text{test}}|^{-1}$, rendering traditional benchmarks outdated for models trained on web-scale data. Data privacy and commercial concerns further complicate contamination issues. Second, traditional static methods fall short in preventing contamination, prompting the creation of dynamic benchmarks. However, our study identifies a lack of standardized criteria for evaluating these dynamic benchmarks. We hope our proposed criteria will offer valuable insights and guide the development of more effective benchmarks.
Current Challenges. Static benchmarking methods face challenges due to a lack of transparency (e.g., label protection) and high assumptions about contaminated models (e.g., post-hoc detection). Although dynamic benchmarks address these limitations, they introduce new issues, such as balancing correctness with scalability. We also observed that some dynamic benchmarks neglect complexity control, resulting in inefficiencies in evaluation.
Future Directions. Dynamic benchmarking offers a new approach for evaluating LLMs, but our study reveals a lack of standardized criteria. Future efforts should focus on standardizing these criteria for dynamic benchmarks.
6 Conclusion
This survey reviews the literature on data contamination in LLM benchmarking, analyzing both static and dynamic approaches. We find that static methods, though consistent, become more vulnerable to contamination as training datasets grow. While dynamic approaches show promise, they face challenges in reliability and reproducibility. Future research should focus on standardized dynamic evaluation, and practical mitigation tools.
Limitations
@@ 6 Conclusion
While this survey provides a comprehensive overview of static and dynamic benchmarking methods for LLMs, there are several limitations to consider. First, due to the rapidly evolving nature of LLM development and benchmarking techniques, some recent methods or tools may not have been fully covered. As benchmarking practices are still emerging, the methods discussed may not yet account for all potential challenges or innovations in the field. Additionally, our proposed criteria for dynamic benchmarking are a first step and may need further refinement and validation in real-world applications. Lastly, this survey focuses primarily on high-level concepts and may not delve into all the fine-grained technical details of specific methods, which may limit its applicability to practitioners seeking in-depth implementation guidelines.

### v2

@@ 4.1 Problem Formulation
4.2 Criteria Summarization and Abstraction
While many dynamic benchmarking methods have been proposed to evaluate LLMs, the criteria for evaluating these benchmarks themselves remain non-standardized. To address this gap, we analyze existing evaluation practices and abstract them into a unified framework. We review over 50 dynamic benchmarking papers, focusing specifically on how they evaluate their own benchmarks. Although many of these works include some form of self-evaluation, the dynamic benchmarking methods are often incomplete, or lack depth. For example, DyVal2 evaluates benchmark complexity and correctness, but does not address the interpretability of the benchmark construction process.
To systematize this landscape, we identify a unified set of evaluation criteria and present them in Table 2. We then assess whether each dynamic benchmark fully supports, partially supports, or does not support each criterion. For instance, in the case of correctness: benchmarks with built-in guarantees—such as those using temporal cutoffs or rule-based generation—are marked as “supported”. Benchmarks generated using LLMs are marked as “partially supported” if they include validation (e.g., human or automated checks); otherwise, they are labeled “not supported”. More guidance for classifying each dynamic benchmark could be found in Appendix D.
Dynamic Mechanisms Benchmark Name Evaluation Criteria
Correctness Scalability Collision Stability of Complexity Diversity Interpretability
@@ 4.2 Criteria Summarization and Abstraction
Self-Evolving (Wang et al., 2025)
@@
4.3 Summarized Evaluation Criteria
@@ 4.3.1 Correctness
The first criterion for evaluating the quality of dynamic benchmarking is Correctness. If the correctness of the generated dataset cannot be guaranteed, the benchmark may provide a false sense of reliability when applied to benchmarking LLMs, leading to misleading evaluations. We quantify the correctness of dynamic benchmarks as
@@
where $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$ represent the input and output of the $i^{th}$ transformation, respectively. The function $\mathcal{G}(\cdot)$ is an oracle that returns the ground truth of its input, ensuring an objective reference for correctness evaluation. For example, the function $\mathcal{G}(\cdot)$ could be a domain-specific annotator. This equation can be interpreted as the expected alignment between the outputs of the transformed data set and their corresponding ground truth values, measured using the scoring function $\mathcal{S}(\cdot)$. A higher correctness score indicates that the dynamic benchmark maintains correctness to the ground truth.
@@ 4.3.2 Scalability
The next evaluation criterion is scalability, which measures the ability of dynamic benchmarking methods to generate large-scale benchmark datasets. A smaller dataset can introduce more statistical errors during the benchmarking process. Therefore, an optimal dynamic benchmark should generate a larger dataset while minimizing associated costs. The scalability of a dynamic benchmark is quantified as
@@
This equation represents the expectation over the entire transformation space, where $\lVert T_{i}(\mathcal{D})\rVert$ is the size of the transformed dataset, and $\lVert\mathcal{D}\rVert$ is the size of the original dataset. The function $\textsf{Cost}(\cdot)$ measures the cost associated with the transformation process, which could include monetary cost, time spent, or manual effort according to the detailed scenarios. This equation could be interpreted as the proportion of data that can be generated per unit cost.
@@ 4.3.3 Collision
One of the main motivations for dynamic benchmarking is to address the challenge of balancing transparent benchmarking with the risk of data contamination. Since the benchmarking algorithm is publicly available, an important concern arises: If these benchmarks are used to train LLMs, can they still reliably reflect the true capabilities of the LLMs? To evaluate the robustness of a dynamic benchmark against this challenge, we introduce the concept of collision in dynamic benchmarking. Collision refers to the extent to which different transformations of the benchmark dataset produce overlapping data, potentially limiting the benchmark’s ability to generate novel and diverse test cases. To quantify collision, we propose the following metrics
$\begin{split}&\textsf{Collision Rate}=\mathbb{E}_{\begin{subarray}{c}i,j=1,\,i\neq j\end{subarray}}^{N}\left[\frac{\lVert\mathcal{D}_{i}\cap\mathcal{D}_{j}\rVert}{\lVert\mathcal{D}\rVert}\right]\\ &\textsf{Repeat}=\mathbb{E}_{i=1}^{N}\left[k\mid k=\min\left\{\bigcup_{j=1}^{k}\mathcal{D}_{j}\supseteq\mathcal{D}_{i}\right\}\right]\end{split}$
Collision Rate measures the percentage of overlap between two independently transformed versions of the benchmark dataset, indicating how much potential contamination among two trials. Repeat Trials quantifies the expected number of transformation trials required to fully regenerate an existing transformed dataset $T_{i}(\mathcal{D})$, providing insight into the benchmark’s ability to produce novel variations. These metrics help assess whether a dynamic benchmark remains effective in evaluating LLM capabilities, even when exposed to potential training data contamination.
4.3.4 Stability of Complexity
Dynamic benchmarks must also account for complexity to help users determine whether a performance drop in an LLM on the transformed dataset is due to potential data contamination or an increase in task complexity. If a dynamic transformation increases the complexity of the seed dataset, a performance drop is expected, even without data contamination. However, accurately measuring the complexity of a benchmark dataset remains a challenging task. Existing work has proposed various complexity metrics, but these are often domain-specific and do not generalize well across different applications. For example, DyVal (Zhu et al., 2024a) proposes applying graph complexity to evaluate the complexity of reasoning problems. Formally, given a complexity measurement function $\Psi(\cdot)$, the stability can be formulated as
@@ 4.3.5 Diversity
The diversity metric can be categorized into two components: external diversity and internal diversity. External diversity measures the variation between the transformed dataset and the seed dataset. Internal diversity quantifies the differences between two transformation trials.
@@
where $\Theta(\cdot)$ is a function that measures the diversity between two datasets. For example, it could be the N-gram metrics or the reference-based metrics, such as BLEU scores.
@@ 4.3.6 Interpretability
Dynamic benchmarking generates large volumes of transformed data, making manual verification costly and challenging. To ensure correctness, the transformation process must be interpretable. Interpretable transformations reduce the need for extensive manual validation, lowering costs. Rule-based or manually crafted transformations are inherently interpretable, while LLM-assisted transformations depend on the model’s transparency and traceability. In such cases, additional mechanisms like explainability tools or human-in-the-loop validation may be needed to ensure reliability and correctness.
@@ 4.4 Existing Work
Table 4 summarizes recent dynamic benchmarks. Dynamic benchmarking methods can be categorized into four types: temporal cutoff, rule-based generation, LLM-based generation, and hybrid.
@@ 4.4.1 Temporal Cutoff
Since LLMs typically have a knowledge cutoff date, using data collected after this cutoff to construct datasets can help evaluate the model while mitigating data contamination. This type of method has been widely adopted to construct reliable benchmarks that prevent contamination Uddin et al. (2024). LiveBench (White et al., 2024) collects questions based on the latest information source, e.g., math competitions from the past 12 months, with new questions added and updated every few months. AntiLeak-Bench (Wu et al., 2025) generates queries about newly emerged knowledge that was unknown before the model’s knowledge cutoff date to eliminate potential data contamination. AcademicEval (Zhang et al., 2024a) designs academic writing tasks on latest arXiv papers. LiveCodeBench (Jain et al., 2024) continuously collects new human-written coding problems from online coding competition platforms like LeetCode. LiveAoPSBench (Mahdavi et al., 2025) collects live math problems from the Art of Problem Solving forum. Forecastbench (Karger et al., 2024) updates new forecasting questions on a daily basis from different data sources, e.g., prediction markets.
Limitations. The collection process typically requires significant human effort (White et al., 2024; Jain et al., 2024), and continuous updates demand ongoing human involvement. Despite the popularity of temporal cutoffs, using recent information from competitions to evaluate LLMs can still lead to data contamination, as these problems are likely to be reused in future competitions (Wu et al., 2025). Verification is often overlooked in these live benchmarks (White et al., 2024).
4.4.2 Rule-Based Generation
The method of rule-based generation synthesizes new test cases based on predefined rules, featuring an extremely low collision probability (Zhu et al., 2024a).
Template-Based. GSM-Symbolic (Mirzadeh et al., 2025) creates dynamic math benchmarks by using query templates with placeholder variables, which are randomly filled to generate diverse problem instances. Mathador-LM (Kurtic et al., 2024) generates evaluation queries by adhering to the rules of Mathador games (Puma et al., 2023) and varying input numbers. MMLU-CF (Zhao et al., 2025) follows the template of multiple-choice questions and generates novel samples by shuffling answer choices and randomly replacing incorrect options with “None of the other choices”.
Table-Based. S3Eval (Lei et al., 2024) evaluates the reasoning ability of LLMs by assessing their accuracy in executing random SQL queries on randomly generated SQL tables.
Graph-Based. In this category, LLMs are evaluated with randomly generated graphs. For instance, DyVal (Zhu et al., 2024a) assesses the reasoning capabilities of LLMs using randomly generated directed acyclic graphs (DAGs). The framework first constructs DAGs with varying numbers of nodes and edges to control task difficulty. For example, in arithmetic reasoning tasks, leaf nodes represent random numeric values, while edges correspond to randomly assigned arithmetic operators. These DAGs are then transformed into natural language descriptions through rule-based conversion. Finally, the LLM is evaluated by querying it for the value of the root node. Similarly, NPHardEval (Fan et al., 2024) evaluates the reasoning ability of LLMs on well-known P and NP problems, such as the Traveling Salesman Problem (TSP). Random graphs of varying sizes are synthesized as inputs for TSP to assess the LLM’s performance. Xie et al. (2024) automatically construct Knights and Knaves puzzles with random reasoning graph.
Limitations. The pre-defined rules may limit sample diversity, and publicly available rule-generated data may increase the risk of in-distribution contamination during training (Tu et al., 2024).
4.4.3 LLM-Based Generation
Benchmark Rewriting. In this category, LLMs are employed to rewrite samples from existing static benchmarks, which may be contaminated. Auto-Dataset (Ying et al., 2024) prompts LLMs to generate two types of new samples: one that retains the stylistics and essential knowledge of the original, and the other that presents related questions at different cognitive levels (Bloom et al., 1956). StructEval (Cao et al., 2024) expands on examined concepts from the original benchmark by using LLMs and knowledge graphs to develop a series of extended questions. ITD (Zhu et al., 2024c) utilizes a contamination detector (Shi et al., 2024) to identify contaminated samples in static benchmarks and then prompts an LLM to rewrite them while preserving their difficulty levels. VarBench (Qian et al., 2024) prompts LLMs to generate new ones.
Interactive Evaluation. In this category, inspired by the human interview process, LLMs are evaluated through multi-round interactions with an LLM  (Li et al., 2023a). LLM-as-an-Interviewer (Kim et al., 2024) employs an interviewer LLM that first paraphrases queries from existing static benchmarks and then conducts a multi-turn evaluation by posing follow-up questions or providing feedback on the examined LLM’s responses. TreeEval (Li et al., 2024c) begins by generating an initial question on a given topic using an LLM. Based on the previous topic and the examined LLM’s response, it then generates follow-up subtopics and corresponding questions to further assess the model. KIEval (Yu et al., 2024) generates follow-up questions based on the evaluated model’s response to an initial question from a static benchmark.
Multi-Agent Evaluation. Inspired by the recent success of multi-agent systems (Guo et al., 2024), multi-agent collaborations are used to construct dynamic benchmarks. Benchmark Self-Evolving (Wang et al., 2025) employs a multi-agent framework to dynamically extend existing static benchmarks, showcasing the potential of agent-based methods. Given a task description, BENCHAGENTS (Butt et al., 2024) leverages a multi-agent framework for automated benchmark creation. It splits the process into planning, generation, verification, and evaluation—each handled by a specialized LLM agent. This coordinated method, with human-in-the-loop feedback, yields scalable, diverse, and high-quality benchmarks.
Limitations. The quality of LLM-generated samples is often uncertain. For instance, human annotation in LatestEval (Li et al., 2023c) reveals that  10% of samples lack faithfulness or answerability. In interactive settings, reliability further depends on the interviewer LLM.
4.4.4 Hybrid Generation
LatestEval (Li et al., 2023c) combines temporal cutoff and LLM-based generation to automatically generate reading comprehension datasets using LLMs on real-time content from sources such as BBC. DARG (Zhang et al., 2024b) integrates LLM-based and graph-based generation. It first extracts reasoning graphs from existing benchmarks and then perturbs them into new samples using predefined rules. C2LEVA (Li et al., 2025) incorporates all three contamination-free construction methods to build a contamination-free bilingual evaluation. TrustGen (Huang et al., 2025) is the first dynamic benchmarking to evaluate trustworthiness across multiple dimensions and model types, including text-to-image, large language, and vision-language models.
5 Discussions
Current Challenges. Benchmarking LLMs is essential for evaluating model performance, but traditional static benchmarks risk data contamination. Dynamic benchmarks address this risk by updating or regenerating test data, aiming to maintain integrity. However, current dynamic methods often lack standardized evaluation criteria, suffer from limited scalability, and offer little interpretability. Many also fail to systematically assess trade-offs like computational overhead and robustness.
Future Directions. Future work should establish standardized evaluation frameworks with criteria such as correctness, diversity, and scalability. Contamination-resilient benchmarks—using temporal filtering, synthetic data, or rule-based generation—can further improve reliability. Dynamic benchmarks should also support continual updates, cross-model applicability, and human-in-the-loop validation. Public update logs and improved interpretability will enhance transparency and trust in LLM evaluation. Future directions also include extending dynamic benchmarking to multi-modal LLMs (Chen et al., 2024b, c).
6 Conclusion
This survey reviews the literature on data contamination in LLM benchmarking, analyzing both static and dynamic methods. We find that static methods, though consistent, become more vulnerable to contamination as training datasets grow. While dynamic methods show promise, they face challenges in reliability and reproducibility. Future research should focus on standardized dynamic evaluation, and practical mitigation tools.
Acknowledgements
Simin and Baishakhi are partially supported in part by CCF 2313055, CCF 2107405, CAREER 2025082, and FAI: 2040961. Dezhi and Tao are partially supported by National Natural Science Foundation of China under Grant No. 623B2006, and Grant No. 92464301. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors.
@@ Acknowledgements
While this survey provides a comprehensive overview of static and dynamic benchmarking methods for LLMs, there are several limitations to consider. First, due to the rapidly evolving nature of LLM development and benchmarking methods, some recent methods or tools may not have been fully covered. As benchmarking practices are still emerging, the methods discussed may not yet account for all potential challenges or innovations in the field. Additionally, our proposed criteria for dynamic benchmarking are a first step and may need further refinement and validation in real-world applications. Finally, this survey focuses primarily on high-level concepts and may not delve into all the fine-grained technical details of specific methods, which may limit its applicability to practitioners seeking in-depth implementation guidelines.